# CounterQuill: Investigating the Potential of Human-AI Collaboration in Online Counterspeech Writing

### v1

@@ Abstract.
Online hate speech has become increasingly prevalent on social media platforms, causing harm to individuals and society. While efforts have been made to combat this issue through content moderation, the potential of user-driven counterspeech as an alternative solution remains underexplored. Existing counterspeech methods often face challenges such as fear of retaliation and skill-related barriers. To address these challenges, we introduce CounterQuill, an AI-mediated system that assists users in composing effective and empathetic counterspeech. CounterQuill provides a three-step process: (1) a learning session to help users understand hate speech and counterspeech; (2) a brainstorming session that guides users in identifying key elements of hate speech and exploring counterspeech strategies; and (3) a co-writing session that enables users to draft and refine their counterspeech with CounterQuill. We conducted a within-subjects user study with 20 participants to evaluate CounterQuill in comparison to ChatGPT. Results show that CounterQuill’s guidance and collaborative writing process provided users a stronger sense of ownership over their co-authored counterspeech. Users perceived CounterQuill as a writing partner and thus were more willing to post the co-written counterspeech online compared to the one written with ChatGPT.
@@
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
Figure 1. The image showcases CounterQuill’s key features: (A) Learning Session (left) - Educates users about hate speech, counterspeech, and their impact through interactive lessons and quizzes. (B) Brainstorming Session (middle) - Guides users to identify key elements of hate speech, such as targeted identities and dehumanizing actions, and collaboratively brainstorm counterspeech strategies. (C) Writing Session (right) - Provides an AI-assisted writing environment with grammar checking, empathetic tone polishing, and personalized suggestions to help users write counterspeech.
@@
The rapid growth of online communication platforms has led to an increase in the prevalence of hate speech, causing significant harm to individuals and society (Holt and Bossler, 2020). A study by the Pew Research Center found that 41% of U.S. adults have experienced online hate speech, with 18% subjected to severe forms such as physical threats, persistent harassment, or sexual harassment (DUGGAN, 2017). Youth and young adults are even more likely to be exposed to online hate. One cross-national study of online hate in Finland, France, Poland, Spain, the United Kingdom, and the United States found that approximately 70% of youth and young adults ages 18 to 25 had seen online hate materials. Exposure rates ranged from 65% in France to almost 79% in Finland, with 73% of Americans being exposed (Reichelmann et al., 2021). To reduce online hate, technology companies have employed human moderators and automated moderation systems. However, prior research suggests that these efforts, particularly the practice of deplatforming – banning users who express hateful views – may not effectively eliminate hate speech and often drives users to less monitored online environments (Horta Ribeiro et al., 2023). The limitations of traditional moderation methods have prompted scholars to explore the potential of user-driven counterspeech as an alternative solution to combat online hate.
Counterspeech is “communication that tries to counteract potential harm brought about by other speech” (Cepollaro et al., 2023; Rieger et al., 2018). With respect to hate speech, counterspeech is a direct response to derogatory or harmful content, intended to undermine or refute the hateful message with the intention of stopping it, reducing its impact, or providing support to those it targeted (Rieger et al., 2018). However, individuals often face significant challenges and barriers when engaging in this practice, such as fear of retaliation as well as emotional and skill-related barriers (Mun et al., 2024; Ping et al., 2024). Previous research has identified such barriers as crucial factors deterring people from engaging in counterspeech (Ping et al., 2024). Individuals with these barriers were found to be less satisfied with their own counterspeech, experienced more difficulty in writing it, and were less likely to perceive their counterspeech as effective (Ping et al., 2024).
Given the barriers that individuals face when engaging in counterspeech, the rapid advancements in Natural Language Processing (NLP) and Large Language Models (LLMs) offer avenues for supporting individuals in combating online hate (Yu et al., 2022; Zhu and Bhat, 2021; Saha et al., 2022; Garland et al., 2020). Researchers have been exploring various aspects of NLP techniques and LLMs, such as context-based hate speech recognition (Yu et al., 2022; Garland et al., 2020) and counterspeech generation (Zhu and Bhat, 2021; Saha et al., 2022). However, despite such potential, LLM-generated counterspeech have several limitations. For example, some worry that AI-generated counterspeech might lack personal touch(Baider, 2023; Ping et al., 2024; Mun et al., 2024). Prior studies show that some also doubt AI’s ability to mimic human emotions such as empathy, which is important when engaging in sensitive topics and countering hate speech (Ping et al., 2024; Hangartner et al., 2021). Furthermore, previous research discovered that counterspeech generated by AI (such as ChatGPT) prevented users from feeling a sense of ownership over the words and ideas expressed because ChatGPT automates the generation without engaging people in the process of writing counterspeech (Kumar, 2024).
To address these challenges, we introduce CounterQuill, an AI-mediated counterspeech writing partner that assists users in composing counterspeech by incorporating their own experiences and understanding. The system offers a co-writing feature that provides writing assistance based on the user’s ideas, allowing them to explore different strategies and tones, such as an empathetic approach. Throughout the interactive process, CounterQuill offers targeted feedback on each user’s counterspeech idea, refining and rewriting the text based on the user’s focus and preferences. By engaging with the system’s feedback, users discover that a user-centric approach — combining AI-provided advice with human-written personal experiences — creates a tailored and personally relevant counterspeech.
To evaluate CounterQuill, we conducted a within-subjects study with 20 participants. Each participant was shown online hate posts and tasked to write counterspeech under two different conditions: (1) using ChatGPT, and (2) using CounterQuill. To minimize potential biases, we randomly assigned the order of conditions: with half the participants starting with our system and the other half beginning with ChatGPT. Specifically, we address the following research questions in our user study to better understand the role of AI in assisting users learn, brainstorm, and draft counterspeech: (RQ1) How does the learning session influence users’ understanding of hate speech, as well as the strategies and empathy-based approaches for counterspeech? (RQ2) How does the brainstorming session guide users through the process of identifying and responding to identity-based dehumanization in online hate speech impact users’ confidence, and perceived ability to engage in counterspeech? (RQ3) How does AI assistance in drafting and refining counterspeech impact users’ perceived self-efficacy in counterspeech writing and their perspectives towards the counterspeech co-written with AI?
@@
Contributions: The intellectual merits of this work lies in the design of an AI-mediated counterspeech system and the potential social impact of human-AI collaboration.
@@
  • We show that CounterQuill assists users in identifying the core elements of hate speech, such as the targeted individual or group’s identity and the dehumanizing actions. The system’s highlighting practice and AI-generated feedback enhance participants’ ability to recognize these key components.
@@
  • We find that CounterQuill increases users’ confidence in crafting empathetic counterspeech. The system’s brainstorming session combined with the co-writing session that provides targeted suggestions and rewriting assistance, empowers users to compose empathetic counter responses.
@@
  • We uncover a relationship between user cognitive effort and behavioral outcomes in AI-mediated communication (AIMC) systems. Our findings suggest that when users invest more time and effort in collaborating with an AI system to craft counterspeech, they develop a stronger sense of ownership over the co-authored counterspeech, increasing their likelihood of posting it on social media.
  • Finally, our system demonstrates broad adaptability to diverse user groups, as it can assist different individuals in countering hate speech. By providing customized guidance and support tailored to users’ unique experiences, attitudes, and skill levels, CounterQuill showcases the potential for AI-mediated tools to help users combat online hate.
@@ 2.4. Ownership in Human-AI Collaboration
Previous research has demonstrated that ownership extends beyond physical or tangible objects, encompassing everything from people to text to ideas (Heider, 1958; Mieczkowski, 2022). Regarding how interaction with AI systems affects the human sense of ownership, past studies have found that the more people engage in their interactions with AI, the more likely they are to develop a strong sense of ownership over the outcomes of their co-creation with the AI (Louie et al., 2021; Pierce et al., 2003; Louie et al., 2020a). For instance, in Louie et al.’s (2021) study of AI music co-creation, participants indicated that they felt the final product did not belong to them due to their minimal control over the AI-generated suggestions (Louie et al., 2021). Furthermore, Draxler et al.’s study of AI-assisted writing suggested that users’ perceived control over AI-generated content and their feeling of leadership in the interaction significantly impact their sense of ownership (Draxler et al., 2024).
@@
Overall, the growing collaboration between humans and AI in content creation has led to an interplay of user effort (Mieczkowski, 2022; Louie et al., 2021), perceived ownership (Mieczkowski, 2022; Kumar, 2024; Louie et al., 2020a; Pierce et al., 2003), and ethical considerations (Kumar, 2024; Ping et al., 2024). Our research aims to further explore the relationship between people’s effort, perceived ownership, and future behavioral outcomes in human-AI collaboration.
3. BACKGROUND AND DESIGN GOALS
We take a human-centered approach in the design of CounterQuill as an interactive human-AI system that empowers users with relevant knowledge (DG1), ideation space (DG2), and collaborative writing assistance (DG3) rather than simply generating counterspeech. By doing so, our system aims to help users engage in the process of crafting effective and empathetic counterspeech that incorporates their preferences and thought-process in countering online hate, rather than soleley delivering a machine-generated output. Our system is designed to achieve the following design goals:
@@ 3.1. Help Users Learn About Hate Speech & Counterspeech and Their Impact (DG1)
3.1.1. Understanding Hate Speech and its Impact
Previous research found that many internet users, especially younger generations, may not fully understand what constitutes hate speech or its harmful effects (Räsänen et al., 2016; Anti-Defamation League, 2019). A survey conducted by the Anti-Defamation League revealed that 65% of American youth were exposed to online hate speech, but only 49% of them recognized and reported hateful content to social media companies (Anti-Defamation League, 2019). This lack of awareness underscores the need for educational initiatives to help users identify hate speech effectively. To address this issue, NGOs have launched campaigns aimed at raising awareness about hate speech and teaching people how to recognize it (Laville, 2016; Schieb and Preuss, 2016).
Previous research has found that hate speech can be expressed both explicitly and implicitly (ElSherief et al., 2021; Gunturi et al., 2024). Explicit hate speech directly attacks social groups or individuals through overt language, such as profanity (Reichelmann et al., 2021; Howard, 2021). In contrast, implicit hate speech is more subtle (ElSherief et al., 2021), often relying on coded language (ElSherief et al., 2021; Ocampo et al., 2023), neologisms (Retta, 2023), sarcasm (Ocampo et al., 2023), or microaggressions (Gunturi et al., 2024) to convey prejudice or discrimination. Thus, same words or phrases can have different meanings depending on the specific situation, the speaker’s intent, and the audience’s perception (Gunturi et al., 2024). For such reasons, it can be particularly difficult for humans to recognize implicit hate speech.
3.1.2. Understanding counterspeech and its Impact
Furthermore, previous research found that the lack of clear guidelines and practices for crafting effective counterspeech is another factor contributing to people’s reluctance to respond to hate speech (Ping et al., 2024). Baider (2020) observed that individuals often struggle to determine what constitutes impactful counterspeech and how to formulate responses that can make a difference (Baider, 2020).
Research has highlighted the potential for unintended consequences when counterspeech is not delivered carefully (Howard, 2021). As Howard’s (2019) study found, counterspeech perceived as aggressive or preachy can sometimes backfire, leading to increased hate speech (Howard, 2021). Humorous counterspeech also has mixed results on hate speech. While humour can sometimes defuse hate speech, attempts at humorous counterspeech in other cases can backfire (Schieb and Preuss, 2016). Compared to other strategies like humour or warning of consequences, Hangartner et al. show that empathy-based counterspeech - counterspeech that humanizes the victim and reminds the perpetrator that people can be hurt by their behaviour - is significantly more likely to lead to apologies or deletion of hateful posts by bad actors (Hangartner et al., 2021). However, most people are not only unaware of the different counterspeech strategies and their potential pitfalls and risks, but which ones may be more effective (Ping et al., 2024). To address these issues, our system is designed to help people learn more about hate speech, counterspeech and their impact.
@@ 3.2. Help Users Ideate Counterspeech Ideas Through a Brainstorming Process (DG2)
Previous research has found that while many people recognize the harmful effects of hate speech, they may feel uninformed or unable to engage in counterspeech effectively. Studies by Mun et al. (2024), Buerger et al. (2021), and Ping et al. (2024) revealed that due to cognitive load, users often hesitate or struggle to create effective counterspeech (Mun et al., 2024; Buerger, 2021; Ping et al., 2024) . This cognitive load arises from the difficulty in accurately identifying the key hate elements in hate speech, such as aspects of an individual or group’s identity and actions or perceptions that dehumanize that person or group (Waltman, 2018; Harel et al., 2020). Furthermore, many people may lack the skills to effectively counter hate speech (Ping et al., 2024). A study by Costello et al. (2017) found that although young adults (from an online survey, N = 963) recognized the importance of countering online hate speech, they were often uncertain about how to effectively counter the key hate elements of hate speech (Costello et al., 2017).
To address these challenges, we adopted the traditional brainstorming approach (TBS) (Gallupe et al., 1991; Yu-Han and Chun-Ching, 2023) and extended it by drawing on prior HCI work on AI technologies that facilitate brainstorming (Lavrič and Škraba, 2023; Zhang et al., 2023; Yu-Han and Chun-Ching, 2023; Cepolina et al., 2022). Researchers in these studies use LLMs to help people generate (Zhang et al., 2023; Yu-Han and Chun-Ching, 2023), refine (Yu-Han and Chun-Ching, 2023; Geerts et al., 2021), and evaluate ideas (Cepolina et al., 2022). For example, Gero et al. developed Metaphoria, an LLM-powered system that assists writers in brainstorming creative ideas by generating metaphorical associations and encouraging diverse creative outputs (Gero et al., 2022). In addition, Zhang et al. created Visar, a human-AI collaborative writing assistant that helps writers brainstorm and revise hierarchical goals within their argumentative writing context (Zhang et al., 2023).
Lubart (2005) and Chiu et al. (2023) (Lubart, 2005; Yu-Han and Chun-Ching, 2023), identified four key components in HCI brainstorming: (1) AI engaging in communication and guidance with humans during the thinking process; (2) AI providing targeted feedback; (3) AI and humans collaborating to organize brainstorming results; and (4) AI encouraging reflective behavior by integrating human-machine collaboration into the creative process. Built on previous research, we developed two AI agents: [Uncaptioned image] Brainstorming Guidance Agent (BGA), which facilitates communication, guidance, and encouragement, and [Uncaptioned image] Brainstorming Feedback Agent (BFA), which provides feedback and collaborates with users to organize the outputs of the brainstorming process. Overall, our brainstorming session aims to help users ideate counterspeech ideas through a brainstorming process assisted by these AI agents.

### v2

@@ Abstract.
Online hate speech has become increasingly prevalent on social media, causing harm to individuals and society. While content moderation has received considerable attention, user-driven counterspeech remains a less explored yet promising approach. However, many people face fear of retaliation and difficulties in articulating effective responses. We introduce CounterQuill, an AI-mediated system that supports users in composing empathetic counterspeech. CounterQuill’s three-step workflow includes: (1) a learning session on hate speech and counterspeech; (2) a brainstorming session to identify key elements of hateful content and develop strategies; and (3) a co-writing session to refine counterspeech. In a within-subjects user study (N=20) comparing CounterQuill to ChatGPT, participants found CounterQuill’s structured guidance and collaboration more empowering, increasing their willingness to share their co-written counterspeech online. We discuss implications for designing AI-assisted, user-centered tools to mitigate online hate while preserving individual voices and a sense of ownership.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
Figure 1. CounterQuill’s three-stage workflow: (A) Learning Session builds users’ understanding of hate speech, counterspeech, and their broader implications through interactive lessons and quizzes; (B) Brainstorming Session guides users in identifying targeted identities and harmful assumptions, then brainstorming counterspeech strategies; and (C) Co-Writing Sessionoffers an AI-assisted environment for drafting, refining, and personalizing counterspeech with grammar checks, empathetic tone suggestions, and user-driven input.
@@ 1. Introduction
Starting last year and intensifying in the last few weeks under the new U.S. administration, many social media companies have begun lifting restrictions on hate speech (Dodds, 2025). Rather than relying on top-down platform moderation, sites like Facebook are increasingly shifting to a community-driven content moderation strategy, placing users in charge of identifying and addressing potentially false or offensive content—including hate speech (Meta, 2025). This shift coincides with a broader rise in harmful online content: the rapid growth of communication platforms has led to an overall increase in hate speech, causing significant harm to individuals and society (Holt and Bossler, 2020). A study by the Pew Research Center found that 41% of U.S. adults have experienced online hate speech, with 18% subjected to severe forms such as physical threats, persistent harassment, or sexual harassment (DUGGAN, 2017). Youth and young adults are even more vulnerable, with a cross-national study reporting that approximately 70% of individuals aged 18–25 have encountered hateful materials (Reichelmann et al., 2021).
Counterspeech is broadly defined as communication intended to counteract potential harm caused by other speech (Cepollaro et al., 2023; Rieger et al., 2018). In the context of hate speech, counterspeech functions as a direct response—aiming not only to refute harmful content and mitigate its impact, but also to support targeted individuals (Rieger et al., 2018). Despite its promise, engaging in online counterspeech poses significant challenges; individuals often encounter barriers such as fear of retaliation and limited writing skills (Mun et al., 2024; Ping et al., 2024). Research indicates that those facing these challenges tend to be less satisfied with their counterspeech, experience greater difficulty in crafting responses, and doubt the overall effectiveness of their interventions (Ping et al., 2024).
Rapid advances in natural language processing (NLP) and large language models (LLMs) have prompted researchers to examine how AI can be leveraged to help users combat online hate (Yu et al., 2022; Zhu and Bhat, 2021; Saha et al., 2022; Garland et al., 2020). Researchers have explored various NLP applications—from hate speech classification (Yu et al., 2022; Garland et al., 2020) and counterspeech tone recognition (Bonaldi et al., 2024) to automated counterspeech generation (Zhu and Bhat, 2021; Saha et al., 2022). However, studies have revealed significant limitations in fully AI-generated counterspeech, including concerns about lack of personal voice and inability to convey authentic human emotions when addressing sensitive topics (Baider, 2023; Ping et al., 2024; Mun et al., 2024; Hangartner et al., 2021). This highlights the need for a more balanced approach—one that leverages AI’s capabilities while preserving the human voice and emotional authenticity of users in counterspeech creation.
To address this, we developed CounterQuill, an AI-mediated counterspeech writing partner that guides users through a three-step process:
  • (1) a learning session to help users better understand the impact and nature of online hate speech and counterspeech
  • (2) a brainstorming session that supports users in identifying key elements of hate speech and exploring counterspeech strategies
  • (3) a co-writing session that helps users draft and refine their counterspeech based on sessions (1) and (2).
The development of CounterQuill was informed by a preliminary survey (N=458) aimed at understanding user perspectives on the role of AI in counterspeech engagement. Survey results showed that while some users saw potential in AI tools like ChatGPT for writing counterspeech, others were concerned about preserving authenticity and personal expression when using AI for counterspeech writing.
In line with established research practices, we evaluated CounterQuill against ChatGPT—a baseline widely used in studies evaluating AI writing tools (Zhang et al., 2023; Reza et al., 2024; Lin et al., 2024). This comparison allowed us to assess whether CounterQuill’s distinctive features—its integrated learning and brainstorming components alongside its collaborative writing environment—provide meaningful advantages over standard AI assistance. Specifically, we examined whether these additional components help users craft counterspeech they perceive as both effective and authentic to their own voice.
To this end, we conducted a within-subjects study with 20 participants, randomly assigning the order of conditions (half starting with CounterQuill and half with ChatGPT) to minimize potential biases from sequencing effects (Ryan and Morgan, 2007). Our work addresses the following research questions:
  • RQ1. How does AI-mediated learning sessions shape users’ understanding of hate speech and counterspeech?
  • RQ2. How does AI-mediated brainstorming sessions affect users’ confidence and ability to identify key elements of hate speech and counterspeech strategies?
  • RQ3. How does engaging in collaborative co-writing with AI influence users’ self-efficacy in composing counterspeech and their perceptions of the co-authored response?
@@
Contributions: The intellectual merits of this work lie in the design of an AI-mediated counterspeech system and the potential social impact of human-AI collaboration.
@@
  • We show that CounterQuill assists users in identifying the core elements of hate speech, such as the targeted individual or group’s identity and the dehumanizing actions. The highlighting practice and AI-generated feedback enhance participants’ ability to recognize these key components.
@@
  • We find that CounterQuill increases users’ confidence in crafting empathetic counterspeech. The system’s brainstorming session combined with the co-writing session that provides targeted suggestions and rewriting assistance, empowers users to compose empathetic counterspeech.
@@
  • We uncover a relationship between user cognitive effort and behavioral outcomes in AI-mediated communication (AIMC) systems. Our findings suggest that when users invest more time and effort in collaborating with an AI system to write counterspeech, they develop a stronger sense of ownership over the co-authored counterspeech, increasing their likelihood of posting it on social media.
  • Finally, we offer key design implications for AI-assisted counterspeech writing that balance system guidance with user autonomy, prioritizing meaningful user engagement with AI to enhance both the quality of Counterspeech and users’ sense of ownership over the co-created content.
@@ 2.4. Ownership in Human-AI Collaboration
Previous research has demonstrated that ownership extends beyond physical or tangible objects, encompassing everything from people to text to ideas (Heider, 1958; Mieczkowski, 2022). Regarding how interaction with AI systems affects the human sense of ownership, past studies have found that the more people engage in their interactions with AI, the more likely they are to develop a strong sense of ownership over the outcomes of their co-creation with the AI (Louie et al., 2021; Pierce et al., 2003; Louie et al., 2020). For instance, in Louie et al.’s (2021) study of AI music co-creation, participants indicated that they felt the final product did not belong to them due to their minimal control over the AI-generated suggestions (Louie et al., 2021). Furthermore, Draxler et al.’s study of AI-assisted writing suggested that users’ perceived control over AI-generated content and their feeling of leadership in the interaction significantly impact their sense of ownership (Draxler et al., 2024).
@@
Overall, the growing collaboration between humans and AI in content creation has led to an interplay of user effort (Mieczkowski, 2022; Louie et al., 2021), perceived ownership (Mieczkowski, 2022; Kumar, 2024; Louie et al., 2020; Pierce et al., 2003), and ethical considerations (Kumar, 2024; Ping et al., 2024). Our research aims to further explore the relationship between people’s effort, perceived ownership, and future behavioral outcomes in human-AI collaboration.
3. DESIGN GOALS
We adopt a human-centered approach in designing CounterQuill as an interactive human-AI system that empowers users with relevant knowledge (DG1), ideation space (DG2), and collaborative writing assistance (DG3) rather than simply generating counterspeech. By doing so, our system aims to help users engage in the process of crafting effective and empathetic counterspeech that incorporates their preferences and thought process in countering online hate, rather than solely delivering a machine-generated output. Our system is designed to achieve the following design goals:
@@ 3.1. Help Users Learn About Hate Speech & Counterspeech and Their Impact (DG1)
Studies show many internet users, especially young people, encounter hate speech regularly but struggle to identify or report it (Räsänen et al., 2016; Anti-Defamation League, 2019). A survey found 65% of American youth had seen online hate speech, yet only half recognized it as harmful—highlighting a critical need for better education on identifying these behaviors (Anti-Defamation League, 2019). To address this issue, organizations have launched campaigns to teach people how to recognize hate speech, but challenges remain due to its complex nature. Hate speech isn’t always obvious. While some attacks use direct slurs or threats (explicit hate speech) (Reichelmann et al., 2021; Howard, 2021), others hide behind sarcasm (Ocampo et al., 2023), coded language (ElSherief et al., 2021; Ocampo et al., 2023), or subtle stereotypes (implicit hate speech) (Gunturi et al., 2024). The subtleties of implicit hate speech make it hard for people to detect, as meaning shifts based on intent, audience, and cultural cues (Gunturi et al., 2024).
Compounding this, if counterspeech is not delivered carefully, it can often escalate tensions in online hate speech discussions (Howard, 2021), as aggressive or lecturing responses tend to provoke further hostility (Howard, 2021; Chung et al., 2023). To address this challenge, researchers have explored more effective intervention strategies. For example, Hangartner et al. show that empathy-based counterspeech - humanizes the victim and reminds the perpetrator that people can be hurt by their behavior - is more likely to lead to apologies or deletion of hateful posts by bad actors (Hangartner et al., 2021). However, most people lack awareness of these counterspeech strategies and their relative effectiveness (Ping et al., 2024). To address this knowledge gap, we developed our learning session aiming to educate users about both hate speech and counterspeech and understand the impact of different response approaches.
@@ 3.2. Help Users Ideate Counterspeech Ideas Through a Brainstorming Process (DG2)
Prior studies have shown that many people hesitate to engage in online counterspeech due to the cognitive load (Mun et al., 2024; Buerger, 2021; Ping et al., 2024). This arises from the difficulty in identifying the key hate elements of hate speech, such as aspects of an individual or group’s identity and the dehumanization of that identity (Waltman, 2018; Harel et al., 2020) - a task that is particularly challenging in implicit hate speech. Furthermore, many people may lack the skills to counter hate speech (Ping et al., 2024). For instance, a study by Costello et al. (2017) found that although participants (N = 963) encountered online hate speech, only 3.1% indicated they would respond with a counterspeech, while others were unsure how to respond effectively(Costello et al., 2017).
To help users overcome such barriers and practice counterspeech skills, we propose a brainstorming approach to guide counterspeech ideation process. We adopted the traditional brainstorming approach (TBS) (Gallupe et al., 1991; Yu-Han and Chun-Ching, 2023) and extended it by drawing on prior HCI studies on AI technologies that facilitate brainstorming (Lavrič and Škraba, 2023; Zhang et al., 2023; Yu-Han and Chun-Ching, 2023; Cepolina et al., 2022). For example, researchers in their studies use LLMs to help people generate (Zhang et al., 2023; Yu-Han and Chun-Ching, 2023), refine (Yu-Han and Chun-Ching, 2023; Geerts et al., 2021), and evaluate ideas (Cepolina et al., 2022). Zhang et al. created Visar, a human-AI collaborative writing assistant that helps writers brainstorm and revise hierarchical goals within their argumentative writing context (Zhang et al., 2023). In addition, Lubart (2005) and Chiu et al. (2023) (Lubart, 2005; Yu-Han and Chun-Ching, 2023), identified four key components in HCI brainstorming: (1) AI engaging in communication and guidance with humans during the thinking process; (2) AI providing targeted feedback; (3) AI and humans collaborating to organize brainstorming results; and (4) AI encouraging reflective behavior by integrating human-AI collaboration into the creative process. Built on previous research, we developed our brainstorming session aiming to help users ideate counterspeech strategy.